OpenStudy (anonymous):

limit x->0, tan^2(x)/(x) find the limit, help me!!

OpenStudy (agent0smith):

Have you learned l'hopitals rule?

OpenStudy (anonymous):

yes, derivate?

OpenStudy (agent0smith):

Yeah i'd try that first, since it's indeterminate form...

OpenStudy (anonymous):

I try to do this but I can't find the solution

OpenStudy (agent0smith):

Can you differentiate tan^2x? Do it like (tanx)^2 and just use the chain rule. Derivative of the x is just 1.

OpenStudy (anonymous):

of course but the problem is in the derivative of tan^2(x)

OpenStudy (agent0smith):

(tanx)^2 is the same as tan^2x... if you can differentiate one, you can differentiate the other...

OpenStudy (agent0smith):

\[\Large \frac{ d }{ dx } f(x)^n = n f(x)^{n-1}*f \prime (x)\]

OpenStudy (anonymous):

yes I know the result is 2tan(x)

OpenStudy (agent0smith):

Follow the chain rule i showed above... you're missing something.

OpenStudy (anonymous):

I can't find the mistake, help me please

OpenStudy (agent0smith):

Look at the chain rule... you're missing the derivative of tanx.

OpenStudy (anonymous):

sec^2x ?

OpenStudy (agent0smith):

Yes, that's what you were missing.

OpenStudy (anonymous):

and how I can to evaluate with 0 if I have sec^2x

OpenStudy (agent0smith):

You can evaluate easily. Secant is 1/cosine...

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

i havee secant but sec^2

OpenStudy (agent0smith):

Yep... so you can evaluate 2tanxsec^2x for x=0.

OpenStudy (anonymous):

1/cosx is for the secant not fot sec^2

OpenStudy (agent0smith):

secx = 1/cosx sec^2x = 1/cos^2x

OpenStudy (agent0smith):

because sec^2x = (1/cosx)^2 ... understand...?